OpenStudy (anonymous):

√81x^6 This is what I got can anyone check this for me and explain? √(9x^3)^2√x

OpenStudy (anonymous):

Not sure where the radicals expand out to. sqrt(81x^6)= sqrt(81)*sqrt(x^6) = 9 sqrt(x^6). Now with the sqrt(x^6) we need to be careful. If x < 0 sqrt(x^6) > 0. We need to retain that information. So as with sqrt(x^2) we need abolute value signs. So sqrt(81x^6)=9abs(x^3)

OpenStudy (anonymous):

@Xavier What do you mean 9abs(x^3) abs?

OpenStudy (anonymous):

absolute value abs(x^3) = |x^3|

OpenStudy (anonymous):

@Xavier oooooh alright

OpenStudy (anonymous):

Xavier 9√(x^3)

OpenStudy (anonymous):

I think I went too far with my posted answer